STAR Weekly IST Meeting Minutes 02/02/2015

1. Operations: Xiaozhi
1) Leakage in cooling system fixed by replacing an O-ring in bypassing valve in the pump. Need to follow up with Bob/Gerrit on replacing the O-ring in the other pump once new O-rings are received.

2) Test on IST alarms planned
2. Calibrations: Yaping
1) Geometry table for Run14 updated per Mustafa's request - last Friday
2) Start working on Run14 pre-production data quality check - this Wednesday
3) Need to document the calibration procedure in Run14 - time scale not decided
3. Software: Bingchu
1) Online HFT monitoring updated - need to check it gives the same results as the old setup
2) One hot chip RDO = 6, ARM = 4, APV = 14, was masked out in Run14, and reactivated in summer 2014. Yaping will check the ped/rms of this chip
3) Will work on the documentation of software based on the latest version of Yaping's write-up